OpenStudy (anonymous):

Factor w^2-36 i get 0

OpenStudy (anonymous):

Factor? or do you mean solve?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

(w-6)(w+6) = 0

OpenStudy (anonymous):

If you solve it you go: w^2 = 36 square root both sides and you get w = +6, -6

OpenStudy (anonymous):

okay thank you

OpenStudy (saifoo.khan):

THnks Man!

OpenStudy (anonymous):

would this factor be written out like this b^2-4= ((b-2))